OpenStudy (anonymous):

which of the following statements is (are) true for all the positive integers? 1) 2+2^2+2^3+...+ 2^n+2n+1-2 2)1+3+5+...+ n=n^3

OpenStudy (kinggeorge):

Is there supposed to be an equals sign somewhere in the first one?

OpenStudy (anonymous):

2^2=2^n+1-2

OpenStudy (kinggeorge):

Should it be \[2+2^2+2^3+...+2^n=2^n+1-2?\]

OpenStudy (anonymous):

no here one sec

OpenStudy (kinggeorge):

Thanks. Best way to do this, is look at examples. Are the following statements true?\[2+2^2=2^3-2\]\[2+2^2+2^3=2^4-1\]

OpenStudy (kinggeorge):

The second equation should have a -2 and not -1.

OpenStudy (anonymous):

the first ones true and second one isnt.

OpenStudy (kinggeorge):

If you take the -2 into account the second equation I wrote is actually true as well. This is a good indication that the overall statement may be true.

OpenStudy (anonymous):

ohh well okay thanks man

OpenStudy (kinggeorge):

The second statement is easier. Is the following true?\[1+3+5=5^3\]

OpenStudy (anonymous):

no false

OpenStudy (kinggeorge):

right. So the second statement is definitely false, and the first is probably true.

OpenStudy (anonymous):

dopeee i get it know

OpenStudy (anonymous):

thanks

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (anonymous):

they get harder!!